OpenStudy (anonymous):

One integer is 9 more than another. Their product is 162. Find integers. PLEASE HELP...

OpenStudy (anonymous):

x = y+9 x*y=162 (y+9)y = 162 y^2+9y=162 (y+4.5)^2-20.25=162 (y+4,5) ^2=182.25 y+4,5=sqrt182.25=13.5 y=9 and x equals 18

OpenStudy (campbell_st):

let the 1st number be x the 2nd number is x + 9 the product \[x(x +9) = 162\] solve \[x^2 + 9x - 162= 0\] factorising (x + 18)(x - 9) = 0 x = -18 and 9

OpenStudy (anonymous):

ah, yes, true, forgot the second answer